LOG-IN
Displaying report 1-1 of 1.
MIF (ITF Control)
Print this report.
yutaro.enomoto - 23:43, Friday 30 November 2018 (7224)Get code to link to this report
Fiber fixing cont'd and improved performance of ALS

[Aso, Yokogawa, Sugimoto, Miyazaki, Enomoto]

Today we installed the pillars for the green fiber on PSL table and POP table, and see if the noise performance of ALS would change or not.
The following pictures show how we fixed the pillars and tied the fiber on them.
The point is we wrapped the fiber with clean cloth and tied them to the pillars by using cable ties.
(This is a little bit another story, but we also fixed the fiber on POP table)
 

And then, after restoring the alignment, we locked ALS CARM loop and saw if how the IR resonace is stable.
Here is the result. This plot shows 30 sec long time series of IR transmission of X arm. It seems the stability has been improved.

Besides, we intentionally put the IR frequency slightly off-resonant to get the quasi-linear signal of the frequency stablity of main laser controlled by ALS.
The time is: 120 sec from 2018/11/30 11:48:40 UTC (GPS 1227613738).

Images attached to this report
7224_20181130154105_img5480.jpg 7224_20181130154113_img5481.jpg 7224_20181130154120_img5483.jpg 7224_20181130154126_img5484.jpg 7224_20181130154133_img5486.jpg 7224_20181130154143_img5488.jpg 7224_20181130154312_keepir.png
Comments related to this report
yutaro.enomoto - 13:52, Saturday 01 December 2018 (7230)

> But can you tell if fastening the fibers is really the solution that removed the 0.2-0.5 Hz fluctuations?

Actually no evidence for that. It might be, for example, due to the different conditions of relative alignment of IR and green.

tomotada.akutsu - 01:36, Saturday 01 December 2018 (7227)

Oh, I forgot reporting this; on Nov 30, when I worked at around Xend TMS, I perhaps slightly happend to touch the periscope for IR on the optical table. So if the IR's postion in the CCD would change, that could be one reason. I'm sorry.

kiwamu.izumi - 06:51, Saturday 01 December 2018 (7228)

Cool! But can you tell if fastening the fibers is really the solution that removed the 0.2-0.5 Hz fluctuations?

yutaro.enomoto - 17:27, Saturday 01 December 2018 (7233)

= BruCo for TMSX_IR_PD (X arm transmission of IR) and REFLAIR_A_RF45_I (PDH signal of X arm) =

[Miyazaki, Enomoto]
Using the data we took yesterday, where IR was kept slightly off-resonant, today we ran BruCo for above channels to try to find the noise sources that limit the current IR stablity holded by ALS.
We found PR2_{BF/TM}_DAMP_{L/P}_IN1 have high coherence with them at 0.4-0.5 Hz.

We also found that PSL_MIC has coherence at 50-55 Hz, and PR3_IM_DAMP_L_IN1 at 47-53 Hz.